Answer is provided need step by step solution

7.2.11 Massachusetts residents overall are more wealthy than the average American. Approximately 70% of all children in Massachusetts live at least 200% above the poverty level. If a random sample of 480 Massachusetts residents was selected, what is the probability that between 65% and 75% live at least 200% above the poverty level? (0.9832)
Answer #1

Solution

Given that,

p = 0.70

1 - p = 1 - 0.70 = 0.30

n = 480

= p = 0.70

  [p ( 1 - p ) / n] = [(0.70 * 0.30) / 480 ] = 0.0209

P( 0.65 < < 0.75)

= P[(0.65 - 0.70) /0.0209 < ( - ) / < (0.75 - 0.70) / 0.0209]

= P(-2.39 < z < 2.39)

= P(z < 2.39) - P(z < -2.39)

Using z table,   

= 0.9916 - 0.0084

= 0.9832